OpenStudy (anonymous):

do i have to use u-substitution to integrate e^xsqrt(8 + e^x) ? i dont know if i have to do that or just use product rule. what is general rule of thumb for u-sub?

jimthompson5910 (jim_thompson5910):

Hint: let u = 8+e^x derive to get du/dx = e^x that turns into du = dx*e^x

jimthompson5910 (jim_thompson5910):

so... e^xsqrt(8 + e^x) dx turns into (dx*e^x)sqrt(8 + e^x) which turns into du*sqrt(u)

OpenStudy (anonymous):

ah yeah you can get rid of the other e^x if you do it that way. i have no idea why i have so much trouble with this but thank you!